OpenStudy (secret-ninja):

Guys, please help, I'll give a medal! I have been stuck on this for quite a while! How do the properties of exponents apply to exponential functions?

OpenStudy (anonymous):

\[(e^x)^a = (e)^{ax}\]

OpenStudy (anonymous):

\[(e^x) \cdot (e^y) = (e)^{x+y}\]

OpenStudy (anonymous):

I think there is no difference, have you got any particular expression where you are stuck?

OpenStudy (secret-ninja):

No, I am answering general questions of things I have already learned so I can be prepared before I take an Exam. I don't remember a lot of the things they asked me. :/ I guess you could say I'm pretty much screwed.

OpenStudy (anonymous):

How can I help you more in this topic?

OpenStudy (anonymous):

\[(\frac{e^x}{e^y}) = e^{x-y}\]